OpenStudy (nevillesnetwork):

Medal & Fan Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

alones (alones):

Start from dividing both sides by \(-3\)

OpenStudy (3mar):

@NevillesNetwork

OpenStudy (eliesaab):

\[ −3|2x + 6| = −12 \\ |2x +6| =4\\ 2x+6= \pm 4 \]

OpenStudy (eliesaab):

Can you finish it?